Jessie is three years older than Miguel. Twenty years from now, the sum of their ages is 51. Find their present ages.
Answer by mananth(16946)


Let Miguel's present age be x
Jessie's age now will be x+3
After 20 years
Miguel's age will be x+20
Jessie'sage will be x+3 +20
x+20 +x+3 +20 =51
2x +43=51
2x =51-43
2x =8
x =8/2
x=4 Miguel's age
Jessie's age 4+3 =7 years
x